# Statistical Taylor Expansion111

### v9

@@ 2.7 Low-Order Approximation
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}\simeq$ $\;(\delta x)^{2}+\frac{3}{2}(\delta x)^{4}+\frac{7}{6}(\delta x)^{6}+\frac{5}{8}(\delta x)^{8}+o((\delta x)^{10});$ (2.41)
@@
$\delta^{2}\sin(x)\simeq$ $\;(\delta x)^{2}\cos(x)^{2}-(\delta x)^{4}(\cos(x)^{2}\frac{3}{2}-\frac{1}{2})$
$+(\delta x)^{6}(\cos(x)^{2}\frac{7}{6}-\frac{1}{2})-(\delta x)^{8}(\cos(x)^{2}\frac{5}{8}-\frac{7}{24})+o((\delta x)^{10});$ (2.43)
@@
$\delta^{2}x^{2}\simeq$  $\;4x^{2}(\delta x)^{2}+2(\delta x)^{4};$ (2.45)
@@ 2.8 Convergence
$\frac{\delta^{2}(x\pm\delta x)^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\nu(\kappa)\sum_{n=1}^{+\infty}(P(x)\kappa)^{2n}(\begin{pmatrix}2c\\ 2n\end{pmatrix}-2\begin{pmatrix}c\\ 2n\end{pmatrix});$ (2.49)
@@ 2.9 Dependency Tracing
$\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=$ $\;(\delta x)^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+(\delta x)^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$ (2.54)
$+2(\delta x)^{m+n}\left(\frac{f^{(m)}_{x}f^{(n)}_{x}}{m!\;n!}\eta(m+n)-\frac{f^{(m)}_{x}}{m!}\eta(m)\;\frac{f^{(n)}_{x}}{n!}\eta(n)\right);$
@@ 3.2.2 Stable
The condition for not stable always happens together with the not monotonic condition, but seldom happens by itself.
@@ 10.1 Summary of Statistical Taylor Expansion
Statistical Taylor Expansion suggests that a calculation result such as its convergence depends on input bounding factors, which may be another measurement of data quality.

### v10

@@ 2.6 Low-Order Approximation
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}\simeq$ $\;\sigma^{2}+\frac{3}{2}\sigma^{4}+\frac{7}{6}\sigma^{6}+\frac{5}{8}\sigma^{8}+o((\delta x)^{10});$ (2.36)
$\delta^{2}\sin(x)\simeq$ $\;\sigma^{2}\cos(x)^{2}-(\delta x)^{4}(\cos(x)^{2}\frac{3}{2}-\frac{1}{2})$
$+\sigma^{6}(\cos(x)^{2}\frac{7}{6}-\frac{1}{2})-\sigma^{8}(\cos(x)^{2}\frac{5}{8}-\frac{7}{24})+o((\delta x)^{10});$ (2.37)
@@
$\delta^{2}x^{2}\simeq$  $\;4x^{2}\sigma^{2}+2\sigma^{4};$ (2.40)
@@ 2.7 Convergence
$\frac{\delta^{2}(x\pm\delta x)^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\nu(\kappa)\sum_{n=1}^{+\infty}(P(x)\kappa)^{2n}\frac{\begin{pmatrix}2c\\ 2n\end{pmatrix}}{2n};$ (2.44)
2.8 Bounding Range
[…]
figure The result variance ratio $\delta^{2}f/\widehat{\delta^{2}f}$ (as shown by the y-axis on the left) and the bounding leakage (as shown by the y-axis on the right) for the different bounding range $\kappa$ (as shown by the x-axis) for the selected $f(x)$ (as shown by the legend), in which $\widehat{\delta^{2}f}$ is the corresponding variance for $\kappa=5$.
[…]
figure The ratio of adjusted $\widehat{\delta^{2}f}$ to the original $\widehat{\delta^{2}f}$ (as shown by the y-axis) for the different bounding range $\kappa$ (as shown by the x-axis) for the selected $f(x)$ (as shown by the legend), in which $\widehat{\delta^{2}f}$ is approximated by normalizing the input variance with $\zeta(2,\kappa)$ so that $\delta^{2}x=(\delta x)^{2}$.
Formula (2.4) defines a bounding range $[\mu-\varrho\sigma,\mu+\kappa\sigma]$ for the underlying distribution with the distribution mean $\mu$ and deviation $\sigma$. The probability for the random variable to be outside $[\mu-\varrho\sigma,\mu+\kappa\sigma]$ is defined as the bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
[…]
$x\pm y:\;\;\;\zeta(2,\kappa)=\frac{\zeta(2,\kappa_{x})(\delta x)^{2}+\zeta(2,\kappa_{y})(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.45)
When $1/\kappa<P(x)$, $\delta^{2}\log(x\pm\delta x)$ or $\delta^{2}(x\pm\delta x)^{c}$ diverges. To approximate $\widehat{\delta^{2}f}$ when $1/5<P(x)$, an artificial $\kappa_{a}<5$ satisfying $P(x)^{2}\leq\zeta(2,\kappa_{a})/\kappa_{a}^{2}$ can be used to normalize the input variance $(\delta x)^{2}$ as $(\delta x)^{2}/\zeta(2,\kappa_{a})$, to achieve $\delta^{2}x=(\delta x)^{2}$ perfectly for all $\kappa$. Figure 2.8 shows that the result variances are close to those of $5<\kappa$, although as a first order approximation, the results depend on the actual $f(x)$. Figure 2.8 shows a universal worst $\kappa\simeq 1.75$, below which the tail effect of Gaussian distributions disappears.
2.9 Statistical Bounding
[…]
figure The measured bounding leakage (as shown by the y-axis) for the different measuring bounding range $\kappa_{s}$ (as shown by the x-axis) and sample count $N$ (as shown by the legend).
[…]
figure The measured bounding range $\kappa$ (as shown by the y-axis on the left) and the corresponding measured bounding leakage $\epsilon(\kappa)$ (as shown by the y-axis on the right) for the different sample count $N$ (as shown by the x-axis) when the underlying distribution is either uniform or normal (as shown by the legend, with different measuring bounding range for the normal distribution).
When sampling from a distribution, the sample mean $\overline{x}$ and sample deviation $\delta x$ approaches the distribution mean $\mu$ and distribution deviation $\sigma$ respectively with the sample count $N$ as $\sqrt{N}$ [5], to result in the measured bounding leakage $\epsilon(\kappa,N)$ for $[\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$. Because $\epsilon(\kappa)\neq\epsilon(\kappa,N)$ for limited $N$, let $\epsilon(\kappa)=\epsilon(\kappa_{s},N)$, in which $\kappa_{s}$ is the measuring bonding range, and $\kappa(\kappa_{s},N)$ is the measured bounding range.
$\epsilon(\kappa)=$  $\;1-\xi(\frac{\kappa}{\sqrt{2}});$ (2.46)
$\epsilon(\kappa_{s},N)=$  $\;1-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x-\overline{x}|}{\sqrt{2}})-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x+\overline{x}|}{\sqrt{2}});$ (2.47)
[…]
Generally, $\kappa$ in Formula (2.4) increases with the sample count $N$. An ideal variance assumes that $N=\infty$ so that $\mu=\overline{x}$ and $\sigma=\delta x$, which is the current common approach. In statistical Taylor expansion, $\kappa(N)$ makes the replacement of $\mu=\overline{x}$ and $\sigma=\delta x$ for calculating $\overline{f}$ and $\delta^{2}f$ using Formula (2.6), (2.3), (2.9), and (2.3) respectively to be statistically correct.
@@ 2.10 Dependency Tracing
$\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=$ $\;\sigma^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+\sigma^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$ (2.52)
$+2\sigma^{m+n}\left(\frac{f^{(m)}_{x}f^{(n)}_{x}}{m!\;n!}\eta(m+n)-\frac{f^{(m)}_{x}}{m!}\eta(m)\;\frac{f^{(n)}_{x}}{n!}\eta(n)\right);$
@@ 3.2.2 Stable
The unstable condition happens rarely when the monotonic convergence is very slow, such as beyond the upper bound $\delta x$ for the exponent $-2.5<c\leq-1.8$ in Figure 3.2.1.
@@ 3.4 Tracking Rounding Error
3.5 Imprecise Taylor Coefficients
[…]
figure The result deviation of ’$\sum_{n=0}^{\infty}((-1)^{n}\pm\delta y)y^{n}$ vs $\delta y$ for difference $\delta x$ in the legend, in which $\frac{1}{1+y}=\sum_{n=0}^{\infty}(-1)^{n}y^{n}$ is the Taylor expansion without added uncertainty $\delta y$ to each Taylor coefficients.
In Figure 3.5, the coefficients for Taylor expansion $\frac{1}{1+y}=\sum_{n=0}^{\infty}(-1)^{n}y^{n}$ are added with uncertainty $\delta y$. The result uncertainty of $\sum_{n=0}^{\infty}((-1)^{n}\pm\delta y)y^{n},y\in N(0,\delta x)$ increases approximately linearly with both $\delta x$ and $\delta y$. $1/(1\pm\delta x)$ has an upper bound $\delta x=0.20002$, while with $0.5$ uncertainty in the Taylor coefficients, the upper bound $\delta x=0.20008$. Although variance arithmetic allows Taylor coefficients to be imprecise, this paper limits the Taylor coefficients to precise values.
@@ 10.1 Summary of Statistical Taylor Expansion
Statistical Taylor Expansion suggests that a calculation result such as its convergence depends on input bounding ranges, which may be another measurement of data quality.